OpenStudy (anonymous):

simplify the flowing expression by applying the odd/even identities, cofunctiom identities, and cosin of a sum or difference identities.

OpenStudy (anonymous):

cos(14)cos(16)+sin(-14)cos(74)

OpenStudy (anonymous):

this is the subtraction angle formula for cosine

OpenStudy (anonymous):

no, that was wrong sorry

OpenStudy (anonymous):

\[\sin(-14)=-\sin(14)\] so this is the same as \[\cos(14)\cos(16)-\sin(-4)\cos(74) \]

OpenStudy (anonymous):

so it is the ADDITION angle formula for cosine

OpenStudy (anonymous):

add \(16+74\) then find the cosine of that number

OpenStudy (superdavesuper):

pic of a triangle like below may help...